I am having problems figuring out what the equation is for this problem. Paula has 400 milliliters of a solution that is 80% alcohol. She plans to dilute it so that it is only 50% alcohol. How much water will she have to add in order to do this?
Answer by stanbon(75887)

Paula has 400 milliliters of a solution that is 80% alcohol. She plans to dilute it so that it is only 50% alcohol. How much water will she have to add in order to do this?
------
Equation:
water + water = water
0.20*400 + 1x = 0.50(400+x)
---
Multiply thru by 100 to get:
20*400 + 100x = 50*400 + 50x
50x = 30*400
x = 240 ml (amt of water needed)
====================
Cheers,
Stan H.
============